OpenStudy (richyw):

consider the surface S created by revolving the curve C:\(z=\cos\left(x^2\right),\, 0\leq x\leq\pi/2\) in the xz-plane around the z-axis. a) Find a function of two variables whos graph is S b)Set up a double integral in polar coordinates giving the surface area A using the technique of areas of parallelograms

OpenStudy (turingtest):

I'm not sure about b) you still need help with a) ?

OpenStudy (turingtest):

around z\[S=\int_a^b2\pi yf(x)ds\]where\[ds=\sqrt{1+[f'(x)]^2}dx\]

OpenStudy (anonymous):

This is a parametrization of the surface \[ P(r,t)=\left\{r \cos (t),r \sin (t),\cos \left(r^2\right)\right\},\quad 0\le r\le \pi/2; 0\le t \le 2 \pi \\ \left|\left|\frac{\partial P(r,t)}{\partial r}\times \frac{\partial P(r,t)}{\partial t}\right|\right|=\sqrt{2 r^4+r^2-2 r^4 \cos \left(2 r^2\right)} \\ A=\int_0^{\pi/2}\int_0^{2\pi}\sqrt{2 r^4+r^2-2 r^4 \cos \left(2 r^2\right)} dr dt \]